Alternative Financing Plans

Frey Co. is considering the following alternative financing plans:

Plan 1 Plan 2
Issue 10% bonds (at face value) $1,040,000 $520,000
Issue preferred $1 stock, $10 par 860,000
Issue common stock, $5 par 1,040,000 700,000

Income tax is estimated at 40% of income.

Determine the earnings per share on common stock, assuming that income before bond interest and income tax is $312,000.

Enter answers in dollars and cents, rounding to two decimal places.

Plan 1 $ Earnings per share on common stock
Plan 2 $ Earnings per share on common stock
Answer #1
Plan 1 Plan 2
Income before bond interest and income tax 312000 312000
Less:Bond interest expense 104000 52000
Income before income tax 208000 260000
Less: Income tax expense 40% 83200 104000
Net income 124800 156000
Less: Preferred Dividends 0 86000
Net income for Common Stockholders 124800 70000
Divide by Outstanding common shares 208000 140000
Earnings per share on common stock 0.60 0.50
Workings:
Plan 1 Plan 2
Outstanding common shares =1040000/5 =700000/5